If angle 1 and angle 2 are supplementary angles and angle 1 is 12 more than angle 2, what is the measure of each angle?
Answer by tommyt3rd(5050)

x+x+12=180
2x+12=180
x+6=90
x=84
the angles are 84 and 96 degrees